# Large Language Models for Multilingual Vulnerability Detection: How Far Are We?

### v1

@@ 4.2. RQ2: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the line level?
Table 6. The percentage of correct detection for line-level multilingual vulnerability detection using two representative techniques across the Top 25 most dangerous CWE-IDs in 2023
@@
Figure 8. The number of correct detections based on CVSS severity at the function-level in PLMs and LLMs
Figure 9. The number of correct detections based on CVSS severity at the line-level in PLMs and LLMs
Observation 10: Across all CVSS severity levels, GPT-4o with ITFSP demonstrates superior capability in identifying multilingual vulnerabilities, showing exceptional performance at both function and line-level detection compared to other studied models. For function-level detection, Figure 8 demonstrates that GPT-4o with ITFSP performs better at detecting High, Medium, and Low severity vulnerabilities. However, GPT-4o with ITFSP performs slightly worse than CodeT5P when detecting Critical severity issues. For example, GPT-4o with ITFSP exhibits 630 detections of High severity vulnerability, surpassing 514 by CodeT5P, 438 by Llama 3 with ZSP, 465 by GPT-4o with FSP, and 518 by GPT-4o with ITZSP. GPT-4o with ITFSP exhibits 305 detections of Critical severity vulnerability, which is slightly lower than 321 by CodeT5P. For line-level detection, Figure 9 shows that GPT-4o with ITFSP outperforms all studied PLMs and LLMs at detecting Critical, High, Medium, and Low severity vulnerabilities. For example, GPT-4o with ITFSP exhibits 134 detections of Critical severity vulnerability, significantly surpassing 76 by CodeT5P, 14 by GPT-4o with (ZSP), 117 by GPT-4o with FSP, and 111 by GPT-4o with ITZSP. These findings underscore that GPT-4o with ITFSP consistently delivers superior performance in correctly detecting vulnerabilities at both function-level and line-level granularity across different CVSS severity categories, reflecting the advantage of incorporating instruction tuning and few-shot prompting in enhancing LLMs.
@@
5. Discussion
5.1. How Does the Model Size Affect the LLM’s Strategy?
@@
Table 7. Performance of Larger LLMs in detecting function-level multilingual vulnerability
@@
Techniques Accuracy Recall Precision F1-Score FPR FNR LLMs (Zero-Shot Prompting) Code Llama (7B) 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 Code Llama (70B) 0.5089 0.7360 0.5081 0.6012 0.7210 0.2640 Llama 3 (8B) 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 Llama 3 (70B) 0.4956 0.7733 0.4991 0.6066 0.7855 0.2267 LLMs (Few-Shot Prompting) Code Llama (7B) 0.4763 0.6251 0.4840 0.5456 0.6743 0.3749 Code Llama (70B) 0.4931 0.6133 0.4968 0.5490 0.6286 0.3867 Llama 3 (8B) 0.4299 0.5260 0.4437 0.4814 0.6673 0.4740 Llama 3 (70B) 0.4659 0.6762 0.4781 0.5602 0.7468 0.3238 LLMs (Instruction Tuning + Zero-Shot Prompting) Code Llama (7B) 0.4783 0.5103 0.4824 0.4959 0.5531 0.4897 Code Llama (70B) 0.5035 0.5339 0.5060 0.5196 0.5273 0.4661 Llama 3 (8B) 0.5192 0.5348 0.5215 0.5281 0.4965 0.4652 Llama 3 (70B) 0.5079 0.6300 0.5087 0.5629 0.6157 0.3700 LLMs (Instruction Tuning + Few-Shot Prompting) Code Llama (7B) 0.5094 0.6408 0.5098 0.5678 0.6226 0.3592 Code Llama (70B) 0.4941 0.5564 0.4973 0.5252 0.5690 0.4436 Llama 3 (8B) 0.4946 0.4848 0.4975 0.4911 0.4955 0.5152 Llama 3 (70B) 0.5093 0.5996 0.5104 0.5514 0.5819 0.4004
@@
Table 8. Performance of Larger LLMs in detecting line-level multilingual vulnerability
@@
Techniques Accuracy Recall Precision F1-Score FPR FNR LLMs (Zero-Shot Prompting) Code Llama (7B) 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 Code Llama (70B) 0.9628 0.1321 0.2016 0.1596 0.0144 0.8679 Llama 3 (8B) 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 Llama 3 (70B) 0.9619 0.3092 0.2960 0.3025 0.0202 0.6908 LLMs (Few-Shot Prompting) Code Llama (7B) 0.9666 0.3018 0.3536 0.3257 0.0151 0.6982 Code Llama (70B) 0.9565 0.2385 0.2161 0.2267 0.0237 0.7616 Llama 3 (8B) 0.8980 0.7509 0.1738 0.2822 0.0980 0.2491 Llama 3 (70B) 0.9745 0.5190 0.5237 0.5214 0.0130 0.4810 LLMs (Instruction Tuning + Zero-Shot Prompting) Code Llama (7B) 0.8771 0.5902 0.1234 0.2042 0.1150 0.4098 Code Llama (70B) 0.9565 0.2568 0.2251 0.2400 0.0243 0.7431 Llama 3 (8B) 0.9415 0.3403 0.1819 0.2371 0.0420 0.6597 Llama 3 (70B) 0.9611 0.2573 0.2654 0.2613 0.0195 0.7427 LLMs (Instruction Tuning + Few-Shot Prompting) Code Llama (7B) 0.8688 0.6221 0.1207 0.2022 0.1243 0.3779 Code Llama (70B) 0.9496 0.3918 0.2345 0.2934 0.0351 0.6082 Llama 3 (8B) 0.9220 0.4720 0.1647 0.2442 0.0657 0.5280 Llama 3 (70B) 0.9752 0.5092 0.5378 0.5231 0.0120 0.4908
@@
As shown in Table 7, Code Llama (70B) and Llama 3 (70B) achieve Accuracies of 0.5089 and 0.4956, respectively, in detecting function-level vulnerability using zero-shot prompting. The Code Llama (70B) shows slight improvement compared to Code Llama (7B), while Llama 3 (70B) performs slightly worse than Llama 3 (8B). With instruction tuning and zero-shot prompting, both Code Llama (70B) and Llama 3 (70B) demonstrate similar performance patterns. When using few-shot prompting, Code Llama (70B) and Llama 3 (70B) achieve Accuracies of 0.4931 and 0.4659, exceeding the performance of their smaller counterparts, Code Llama (7B) at 0.4763 and Llama 3 (8B) at 0.4299. When combining instruction tuning with few-shot prompting, Code Llama (70B) and Llama 3 (70B) perform similarly to their smaller versions, achieving Accuracies of 0.4941 and 0.5093 compared to 0.5094 and 0.4946, respectively. As shown in Table 8, for line-level detection, Code Llama (70B) reaches F1-Scores of 0.1596 and 0.2267 when using zero-shot prompting and few-shot prompting, which is degraded from 0.2254 and 0.3257 by Code Llama (7B). Llama 3 (70B) is superior to its smaller version while using zero-shot and few-shot prompting, with F1-Score of 0.3025 and 0.5214 compared to 0.2720 and 0.2822, respectively. Instruction tuning can better improve the larger LLMs at detecting line-level vulnerability. For example, when using instruction tuning with few-shot prompting, Code Llama (70B) and Llama 3 (70B) reach F1-Scores of 0.2934 and 0.5231, significantly superior to 0.2022 by Code Llama (7B) and 0.2442 by Llama 3 (8B).
@@
5.2. Comparison between Reasoning LLMs and Non-Reasoning LLMs
Table 9. Comparison between reasoning and non-reasoning LLMs.
Techniques Accuracy Recall Precision F1-Score FPR FNR Non-reasoning LLMs (Function-level) DeepSeek-Coder 0.5005 0.1796 0.5097 0.2656 0.1748 0.8204 Code Llama 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 Llama 3 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 GPT-3.5-Turbo 0.4827 0.6183 0.4888 0.5459 0.6544 0.3817 GPT-4o 0.4985 0.3631 0.5020 0.4214 0.2632 0.7399 Reasoning LLMs (Function-level) DeepSeek-R1 0.5217 0.6585 0.5193 0.5807 0.6167 0.3415 QwQ-plus 0.5346 0.4838 0.5418 0.5111 0.4141 0.5162 Non-reasoning LLMs (Line-level) DeepSeek-Coder 0.9536 0.3002 0.2246 0.2570 0.0284 0.6997 Code Llama 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 Llama 3 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 GPT-3.5-Turbo 0.9557 0.3215 0.2475 0.2797 0.0229 0.8393 GPT-4o 0.9537 0.4286 0.2700 0.3313 0.0318 0.5714 Reasoning LLMs (Line-level) DeepSeek-R1 0.9692 0.2184 0.3698 0.2746 0.0102 0.7816 QwQ-plus 0.9655 0.2953 0.3344 0.3136 0.0161 0.7047

### v2

@@ 3.2 RQ2: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the line level?
Table 6: The percentage of correct detection for line-level multilingual vulnerability detection using two representative techniques across the Top 25 most dangerous CWE-IDs in 2023
@@
Observation 10: Most vulnerabilities are relatively easy to detect at the function-level but challenging to detect at the line-level. To compare the overlap between vulnerabilities detected at the function-level and line-level, we consider a CWE detectable by the language models only if all 5 techniques in Table 5 and Table 6 can detect it. Otherwise, we consider the CWE undetectable by the language models. For example, CWE-276 (Incorrect Default Permissions) cannot be detected by the language models since the percentage of correct detection is 0 on both GPT-4o (ITZSP) and GPT-4o (ITFSP). In this case, we found 13 CWEs that can be detected at both function-level and line-level: CWE-79, CWE-89, CWE-416, CWE-78, CWE-20, CWE-352, CWE-476, CWE-190, CWE-502, CWE-77, CWE-918, CWE-94, and CWE-863. Additionally, 11 CWEs can be detected at function-level but not at line-level. For example, CWE-787, CWE-125, and CWE-22, three of the most dangerous CWEs, can be effectively detected at function-level but fail to be detected at line-level. Finally, there is 1 exception (CWE-276) that cannot be detected at either function-level or line-level. These results demonstrate that language models exhibit higher detection capabilities at the function-level than at the line-level, with some of the most critical vulnerabilities being identifiable only through function-level analysis.
Figure 8: The number of correct detections based on CVSS severity at the function-level in PLMs and LLMs
Figure 9: The number of correct detections based on CVSS severity at the line-level in PLMs and LLMs
Observation 11: Across all CVSS severity levels, GPT-4o with ITFSP demonstrates superior capability in identifying multilingual vulnerabilities, showing exceptional performance at both function and line-level detection compared to other studied models. For function-level detection, Figure 8 demonstrates that GPT-4o with ITFSP performs better at detecting High, Medium, and Low severity vulnerabilities. However, GPT-4o with ITFSP performs slightly worse than CodeT5P when detecting Critical severity issues. For example, GPT-4o with ITFSP exhibits 630 detections of High severity vulnerability, surpassing 514 by CodeT5P, 438 by Llama 3 with ZSP, 465 by GPT-4o with FSP, and 518 by GPT-4o with ITZSP. GPT-4o with ITFSP exhibits 305 detections of Critical severity vulnerability, which is slightly lower than 321 by CodeT5P. For line-level detection, Figure 9 shows that GPT-4o with ITFSP outperforms all studied PLMs and LLMs at detecting Critical, High, Medium, and Low severity vulnerabilities. For example, GPT-4o with ITFSP exhibits 134 detections of Critical severity vulnerability, significantly surpassing 76 by CodeT5P, 14 by GPT-4o with (ZSP), 117 by GPT-4o with FSP, and 111 by GPT-4o with ITZSP. These findings underscore that GPT-4o with ITFSP consistently delivers superior performance in correctly detecting vulnerabilities at both function-level and line-level granularity across different CVSS severity categories, reflecting the advantage of incorporating instruction tuning and few-shot prompting in enhancing LLMs.
@@
4 Discussion
4.1 A Fine-grained Analysis of Code Structural Properties
Figure 10: Qualitative analysis of code structural properties for samples where language models exhibit divergent performance. The bars represent the median values of five complexity metrics for samples correctly detected only by the LLM (Blue) versus only by the PLM (Orange).
To gain deeper insights into the performance differences between PLMs and LLMs, we conducted a granular qualitative analysis of code structural properties. We utilized CodeT5P as the representative PLM and instruction-tuned GPT-4o (with few-shot prompting) as the representative LLM, since they are the best-performing language model on each category. Our analysis focuses on samples where the models exhibited divergent performance. Specifically, we isolated 608 functions from the test set that were correctly classified only by the LLM (where the PLM failed) and 370 functions correctly classified only by the PLM (where the LLM failed).
To understand “why” behind language model failures and success, we selected five metrics that serve as proxies for the cognitive and structural difficulty of code analysis (Peitek et al., 2021; Weissberg et al., 2025):
  • Cyclomatic Complexity (CC) (McCabe, 1976). Measures the number of linearly independent paths through the code. High CC indicates intricate branching (loops, conditionals), which increases the difficulty of tracking control flow to identify logic-based vulnerabilities.
  • Nesting Degree (Harrison and Magel, 1981). Represents the maximum depth of nested structures (e.g., loops inside loops). Deep nesting obscures variable scope and lifetime, making it harder to detect data-flow anomalies or resource management errors.
  • Token Count & Lines of Code (LOC) (Halstead, 1977) . These measure the volume of the code. In vulnerability detection, longer contexts dilute the model’s attention, making it harder to isolate the specific “needle in the haystack”, the vulnerable line, amidst benign code.
  • Parameter Count (Chidamber and Kemerer, 1994). Indicates the number of arguments in a function signature. High parameter counts often correlate with complex interfaces and state dependencies, complicating the detection of improper input validation or taint propagation.
The structural comparison, summarized in Figure 10, reveals distinct operational profiles for each model architecture when analyzing median values. The most significant differentiator between the models is their ability to handle logical complexity. As evidenced by the Median CC metric, the total median for samples correctly detected only by the LLM is 4.0, compared to a median CC of 2.0 for samples detected only by the PLM. This disparity is particularly stark in languages like C#, where the LLM-exclusive successes had a median CC of 3.0 while the PLM-exclusive successes reached 5.0. Overall, the lower total median for the PLM indicates that CodeT5P struggles to maintain coherence in code characterized by heavy branching, whereas LLM demonstrates superior reasoning capabilities in tracing complex execution paths.
A similar trend is evident when analyzing metrics related to code verbosity and length. The LLM exhibited a clear advantage in processing larger contexts, with the median for its exclusive correct predictions reaching 135.5 tokens and 19.0 lines of code (LOC). Conversely, the samples where the PLM uniquely succeeded were notably shorter, with median values of 96.0 tokens and 15.0 LOC. This data suggests that the PLM’s performance degrades as the input volume increases, likely due to fixed context window constraints or limited attention capacity. LLM, however, remains more robust in verbose scenarios, identifying vulnerabilities where the PLM fails due to information overload.
While the LLM generally dominates in measures of scale and logical complexity, the analysis of Median Nesting Degree offers a nuanced counterpoint. The total median nesting degree for the PLM-exclusive successes was 3.0, which is higher than the 2.0 observed for the LLM’s unique successes. This reversal, notably seen in JavaScript where the PLM handled a median nesting of 4.0 compared to the LLM’s 2.0, implies that while CodeT5P falters with broad logical branching, it remains highly effective at capturing local structural patterns even when they are deeply nested. Regarding interface complexity, the Median Parameter Count also reflects this trend; the LLM correctly classified more complex functions in languages like C# (median of 2.0 parameter count vs 1.0 for PLM), though the total median for both models sat at 2.0.
The data highlights a trade-off between complexity reasoning and pattern recognition. The PLM (CodeT5P) falters primarily when faced with higher logical complexity and extended contexts. In contrast, the LLM (GPT-4o) occasionally fails on samples that are structurally simpler and shorter but may contain specific local patterns that the PLM’s training objective captures more effectively.
4.2 How Does the Model Size Affect the LLM’s Strategy?
@@
Table 7: Performance of Larger LLMs in detecting function-level multilingual vulnerability
@@
Techniques Accuracy Recall Precision F1-Score FPR FNR MCC AUC LLMs (Zero-Shot Prompting) Code Llama (7B) 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 -0.0585 0.4851 Code Llama (70B) 0.5089 0.7360 0.5081 0.6012 0.7210 0.2640 0.0169 0.5075 Llama 3 (8B) 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 0.0010 0.5005 Llama 3 (70B) 0.4956 0.7733 0.4991 0.6066 0.7855 0.2267 -0.0147 0.4939 LLMs (Few-Shot Prompting) Code Llama (7B) 0.4763 0.6251 0.4840 0.5456 0.6743 0.3749 -0.0515 0.4754 Code Llama (70B) 0.4931 0.6133 0.4968 0.5490 0.6286 0.3867 -0.0157 0.4924 Llama 3 (8B) 0.4299 0.5260 0.4437 0.4814 0.6673 0.4740 -0.1440 0.4293 Llama 3 (70B) 0.4659 0.6762 0.4781 0.5602 0.7468 0.3238 -0.0779 0.4647 LLMs (Instruction Tuning + Zero-Shot Prompting) Code Llama (7B) 0.4783 0.5103 0.4824 0.4959 0.5531 0.4897 -0.0429 0.5091 Code Llama (70B) 0.5035 0.5339 0.5060 0.5196 0.5273 0.4661 0.0066 0.5033 Llama 3 (8B) 0.5192 0.5348 0.5215 0.5281 0.4965 0.4652 0.0383 0.5192 Llama 3 (70B) 0.5079 0.6300 0.5087 0.5629 0.6157 0.3700 0.0148 0.5072 LLMs (Instruction Tuning + Few-Shot Prompting) Code Llama (7B) 0.5094 0.6408 0.5098 0.5678 0.6226 0.3592 0.0188 0.5091 Code Llama (70B) 0.4941 0.5564 0.4973 0.5252 0.5690 0.4436 -0.0126 0.4937 Llama 3 (8B) 0.4946 0.4848 0.4975 0.4911 0.4955 0.5152 -0.0107 0.4946 Llama 3 (70B) 0.5093 0.5996 0.5104 0.5514 0.5819 0.4004 0.0180 0.5088
@@
Table 8: Performance of Larger LLMs in detecting line-level multilingual vulnerability
@@
Techniques Accuracy Recall Precision F1-Score FPR FNR MCC AUC LLMs (Zero-Shot Prompting) Code Llama (7B) 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 0.2043 0.6014 Code Llama (70B) 0.9628 0.1321 0.2016 0.1596 0.0144 0.8679 0.1448 0.5589 Llama 3 (8B) 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 0.2730 0.7138 Llama 3 (70B) 0.9619 0.3092 0.2960 0.3025 0.0202 0.6908 0.2830 0.6445 LLMs (Few-Shot Prompting) Code Llama (7B) 0.9666 0.3018 0.3536 0.3257 0.0151 0.6982 0.3097 0.6433 Code Llama (70B) 0.9565 0.2385 0.2161 0.2267 0.0237 0.7616 0.2047 0.6074 Llama 3 (8B) 0.8980 0.7509 0.1738 0.2822 0.0980 0.2491 0.3295 0.8265 Llama 3 (70B) 0.9745 0.5190 0.5237 0.5214 0.0130 0.4810 0.5083 0.7530 LLMs (Instruction Tuning + Zero-Shot Prompting) Code Llama (7B) 0.8771 0.5902 0.1234 0.2042 0.1150 0.4098 0.2296 0.7376 Code Llama (70B) 0.9565 0.2568 0.2251 0.2400 0.0243 0.7431 0.2182 0.6163 Llama 3 (8B) 0.9415 0.3403 0.1819 0.2371 0.0420 0.6597 0.2207 0.6491 Llama 3 (70B) 0.9611 0.2573 0.2654 0.2613 0.0195 0.7427 0.2413 0.6189 LLMs (Instruction Tuning + Few-Shot Prompting) Code Llama (7B) 0.8688 0.6221 0.1207 0.2022 0.1243 0.3779 0.2329 0.7488 Code Llama (70B) 0.9496 0.3918 0.2345 0.2934 0.0351 0.6082 0.2786 0.6784 Llama 3 (8B) 0.9220 0.4720 0.1647 0.2442 0.0657 0.5280 0.2464 0.7031 Llama 3 (70B) 0.9752 0.5092 0.5378 0.5231 0.0120 0.4908 0.5105 0.7485
@@
As shown in Table 7, Code Llama (70B) and Llama 3 (70B) achieve Accuracies of 0.5089 and 0.4956, respectively, in detecting function-level vulnerability using zero-shot prompting. Code Llama (70B) shows slight improvement with a positive MCC of 0.0169 and AUC of 0.5075 compared to Code Llama (7B), while Llama 3 (70B) performs slightly worse than Llama 3 (8B), yielding a negative MCC. With instruction tuning and zero-shot prompting, both Code Llama (70B) and Llama 3 (70B) demonstrate similar performance patterns. When using few-shot prompting, Code Llama (70B) and Llama 3 (70B) achieve Accuracies of 0.4931 and 0.4659, exceeding the performance of their smaller counterparts, Code Llama (7B) at 0.4763 and Llama 3 (8B) at 0.4299. However, the MCC values across few-shot settings generally remain negative, indicating that detecting function-level vulnerabilities remains challenging for these larger LLMs. When combining instruction tuning with few-shot prompting, Code Llama (70B) and Llama 3 (70B) perform similarly to their smaller versions, achieving Accuracies of 0.4941 and 0.5093 compared to 0.5094 and 0.4946, respectively.
As shown in Table 8, for line-level detection, Code Llama (70B) reaches F1-Scores of 0.1596 and 0.2267 when using zero-shot prompting and few-shot prompting, which is degraded from 0.2254 and 0.3257 by Code Llama (7B). This is further corroborated by the MCC and AUC metrics. For example, in zero-shot prompting, Code Llama (70B) drops to an MCC of 0.1448 and AUC of 0.5589, compared to the 7B model’s MCC of 0.2043 and AUC of 0.6014. Llama 3 (70B) is superior to its smaller version while using zero-shot and few-shot prompting, with F1-Score of 0.3025 and 0.5214 compared to 0.2720 and 0.2822, respectively. Notably, in the few-shot setting, Llama 3 (70B) achieves strong performance with an MCC of 0.5083 and an AUC of 0.7530. Instruction tuning can better improve the larger LLMs at detecting line-level vulnerability. For example, when using instruction tuning with few-shot prompting, Code Llama (70B) and Llama 3 (70B) reach F1-Scores of 0.2934 and 0.5231, significantly superior to 0.2022 by Code Llama (7B) and 0.2442 by Llama 3 (8B). This improvement is reflected in the robustness of their discrimination, with Llama 3 (70B) maintaining a relative strong MCC of 0.5105 in this configuration.
@@
4.3 Comparison between Reasoning LLMs and Non-Reasoning LLMs
Table 9: Comparison between reasoning and non-reasoning LLMs.
Techniques Accuracy Recall Precision F1-Score FPR FNR MCC AUC Non-reasoning LLMs (Function-level) DeepSeek-Coder 0.5005 0.1796 0.5097 0.2656 0.1748 0.8204 0.0063 0.5024 Code Llama 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 -0.0585 0.4851 Llama 3 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 0.0010 0.5005 GPT-3.5-Turbo 0.4827 0.6183 0.4888 0.5459 0.6544 0.3817 -0.0375 0.4819 GPT-4o 0.4985 0.3631 0.5020 0.4214 0.2632 0.7399 -0.0035 0.4985 GRACE 0.5061 0.1673 0.5323 0.2546 0.1495 0.8327 0.0244 0.5089 Reasoning LLMs (Function-level) DeepSeek-R1 0.5217 0.6585 0.5193 0.5807 0.6167 0.3415 0.0435 0.5209 QwQ-plus 0.5346 0.4838 0.5418 0.5111 0.4141 0.5162 0.0701 0.5349 Non-reasoning LLMs (Line-level) DeepSeek-Coder 0.9536 0.3002 0.2246 0.2570 0.0284 0.6997 0.2362 0.6359 Code Llama 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 0.2043 0.6014 Llama 3 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 0.2730 0.7138 GPT-3.5-Turbo 0.9557 0.3215 0.2475 0.2797 0.0268 0.6785 0.2596 0.6473 GPT-4o 0.9537 0.4286 0.2700 0.3313 0.0318 0.5714 0.3175 0.6984 Reasoning LLMs (Line-level) DeepSeek-R1 0.9692 0.2184 0.3698 0.2746 0.0102 0.7816 0.2694 0.6041 QwQ-plus 0.9655 0.2953 0.3344 0.3136 0.0161 0.7047 0.2966 0.6396